. A charge of 20 x 10βˆ’6 C is 20 cm from another charge of 20 x 60βˆ’6 C




a. Find the force between them.




b. Determine the 𝐸𝑝 midway between them.




c. How much force will be experienced by an electron placed midway between them?





1
Expert's answer
2022-02-23T12:11:05-0500

Solution

a) force is between them,

"F=\frac{kqq'}{r^2}\\=\frac{9*10^9*20*20*10^{-12}}{0.2*0.2}\\=90N"

b) Determine the 𝐸𝑝 midway between them is zero.

c) force will be experienced by an electron placed midway between them is zero.